# Jack Hong   , Chenxiao Zhao, ChengLIn Zhu, Weiheng Lu, Guohai Xu   , XingYu

### v2

@@ 1 Introduction
These tool-use capabilities can be categorized into two types: (i) Operation tools: Current models cannot perform complex operations on visual or numerical data, including fine-grained image manipulations (e.g., cropping, measuring) and quantitative computations. This limits their capacity to reason about detailed visual content or solve mathematical problems. (ii) Information retrieval tools: Models cannot proactively access up-to-date external knowledge, which often leads to outdated conclusions or statements without verifiable sources. Some recent works attempt to rely on a single tool. For example, as shown in Figure 1 (b), DeepEyes zheng2025deepeyes uses cropping to achieve fine-grained perception, but due to the lack of information retrieval capability, DeepEyes cannot correctly determine the category based solely on its internal knowledge. In contrast, although MMSearch-R1 wu2025mmsearch can perform search, it lacks fine-grained perception, leading to retrieval failures. A substantial gap remains between existing approaches and truly agentic multimodal models. While o3 (o3) has explored “thinking with image” reasoning pattern that combines operations and search, how to realize such capabilities remains unclear.
@@
We first follow the setup of DeepEyes (zheng2025deepeyes) and apply reinforcement learning directly on Qwen2.5-VL (bai2025qwen2), but find that limited inherent tool-use capability prevents stable tool invocation. This highlights the need for a cold-start stage to establish reliable tool-use patterns. Thus, we curate a high-quality dataset that spans diverse scenarios, including perception, reasoning, and search tasks. After cleaning, we apply two filters: (i) difficulty filtering, retaining only questions unsolvable by the base model, and (ii) tool-benefit classification, keeping cases where tool use improves accuracy. Data are split into two subsets: tool-solvable examples for RL and harder unsolved cases for cold start, further augmented with long chain-of-thought trajectories. Supervised fine-tuning on this cold-start dataset enables the model to acquire basic tool-use patterns and deeper reasoning, after which RL further strengthens tool invocation. Notably, we rely only on two simple rewards, accuracy and format, without complex reward engineering (su2025pixel).
@@
Besides, we evaluate DeepEyesV2 on benchmarks covering real-world understanding, mathematical reasoning, and search-intensive tasks. DeepEyesV2 outperforms both general-purpose MLLMs and prior specific reasoning approaches. Specifically, on real-world understanding benchmarks, DeepEyesV2 surpasses even Qwen2.5-VL-32B in some benchmarks through effective tool use. On reasoning tasks, DeepEyesV2 shows preformance gains across multiple benchmarks, including +7.1+7.1 on MathVerse ($52.7\%$ accuracy). On search benchmarks, DeepEyesV2 delivers strong advantages, reaching $63.7\%$ on MMSearch (jiang2024mmsearch), far beyond the MMSearch-R1 (wu2025mmsearch) ($53.8\%$). These results demonstrate that by reliably invoking tools, DeepEyesV2 extends its comprehensive capabilities, achieving accurate and advanced reasoning.
@@ 2 Related Works
Multimodal Large Language Models. The field of multimodal large language models (MLLMs) has witnessed rapid progress in recent years. Early efforts mainly focus on combining pretrained visual encoders with large language models through lightweight adapters or projection layers, enabling basic vision–language alignment and simple multimodal understanding (li2023blip; liu2023visual; liu2024improved; bai2023qwen; chen2024internvl). Subsequently, more powerful architectures such as Qwen2.5-VL (bai2025qwen2), LLaVA-OneVision (li2024llava), and InternVL3 (zhu2025internvl3), expand the training scale and integrated more diverse visual data, significantly improving performance on benchmarks of visual question answering, captioning, and general perception tasks. Recently, some OmniMLLMs (li2025baichuan; zhao2025r1; fu2024vita; jain2024ola; hong2025worldsense) are capable of processing a mix of modalities like speech, video, and images simultaneously. However, existing MLLMs remain largely passive: they can interpret multimodal inputs and generate answers, but lack the ability to actively invoke external tools for computation or knowledge retrieval, which limits their reliability in complex reasoning tasks.
Thinking with Images. The paradigm of “Think with Image” is first introduced by o3 (o3), which demonstrated that multimodal models can interleave reasoning with iterative visual analysis, actively manipulating images to support step-by-step problem solving. Many works attempt to reproduce such capabilities. Most approaches (su2025pixel; lai2025mini; liu2025visual; fan2025grit; jiang2025vlm; zhang2025chain) adopt a two-stage training pipeline, where a cold-start phase is followed by reinforcement learning. In contrast, DeepEyes (zheng2025deepeyes) only adopts reinforcement learning alone and incentivizes the “Think with Image” behaviors, leading to strong reasoning performance. However, the majority of these efforts employ a rather limited tool set, typically restricted to region cropping for fine-grained perception. To improve generality, PyVision (zhao2025pyvision) and Thyme (zhang2025thyme) utilize code execution to enable more flexible visual operations. Despite this progress, these models remain constrained to image manipulation only, and are unable to handle knowledge-intensive questions where access to up-to-date external information is essential.
@@
Search-oriented Reasoning. To mitigate the inherent knowledge limitations of large multimodal language models, a growing line of work explores augmenting them with external knowledge acquisition. Early approaches commonly adopt the retrieval-augmented generation (RAG) paradigm (song2025r1; jin2025search), where relevant information is retrieved from a pre-constructed knowledge base and fed into the model. While effective, this paradigm remains constrained by the static and finite nature of the underlying corpus. To overcome these limitations, more recent studies attempt to leverage online search to dynamically access broader and up-to-date information (zheng2025deepresearcher). Beyond purely textual queries, some efforts extend search into the multimodal domain, enabling retrieval of not only documents but also images, charts, or other media forms relevant to the task (wu2025mmsearch; wang2025vrag). These advances highlight the potential of search-augmented reasoning to complement perception and tool-use capabilities, ultimately broadening the scope of problems that multimodal models can effectively address.
@@ 3.1 Overall Pipeline
Similar to DeepEyes (zheng2025deepeyes), DeepEyesV2 is an agentic multimodal model, but with extended tool-use capabilities beyond simple cropping. In DeepEyesV2, programmatic code execution and web retrieval are treated as complementary and interleavable tools inside a single reasoning trajectory (see Figure 3). Given an image input and the corresponding user query, DeepEyesV2 first generates an initial reasoning plan, and explicitly determines whether this question can be solved directly through internal reasoning or requires tool invocation. If tool use is necessary, DeepEyesV2 emits executable Python code or issues web search queries. Code execution is carried out in a sandboxed environment and can produce structured outputs such as transformed images, numerical measurements, computed arrays, plots, or execution logs. Image queries using the original whole image are submitted via SerpAPI and return the top five visually matched webpages (each with a thumbnail and title). Text queries return the five most relevant webpages, along with titles and snippets. All tool outputs are converted into observations and appended to model’s context. DeepEyesV2 then thinks further in light of these observations and may plan further tool invocations (either additional code, further searches, or both), iterating this reasoning–tool–integration loop until a conclusive answer is produced.
@@ 3.2 Pioneer Experiments
To investigate whether MLLMs can directly acquire tool-use ability through reinforcement learning, we first conduct a pioneer experiment on Qwen2.5-VL (bai2025qwen2) following DeepEyes (zheng2025deepeyes). As shown in Figure 4, during training, we observe that in the early stages model occasionally attempts to produce Python code, but these outputs are often buggy or fail to execute, indicating that existing MLLMs struggle to generate stable and reliable code. As training continues, model gradually abandons code generation and converges to producing only short reasoning chains followed by direct answers, thereby bypassing tool use. Then, to encourage tool invocation, we incorporate the tool usage bonus mechanism from DeepEyes, which explicitly rewards the generation of code. With this additional signal, model is indeed able to produce correct and runnable code in the early stages, suggesting that the mechanism can enforce coding ability. However, with continued training a new degeneration emerges: model’s behavior converged to emitting exactly one code block per query, and this single block typically consists of non-executable, placeholder comments rather than meaningful code, revealing the phenomenon of reward hacking. This pioneer experiment highlights that existing MLLMs cannot reliably learn complex tool use through direct RL alone, motivating the need for a cold start to bootstrap model’s tool invocation ability.
@@ 3.3 Training Data Curation
Data Collection. Pioneer experiments have highlighted the necessity of constructing a high-quality dataset for supervised fine-tuning to explicitly guide model to learn how to generate executable code and perform tool invocations. Following DeepEyes (zheng2025deepeyes), we collect data in accordance with the following principles: (i) Diverse tasks and image distribution. We incorporate varied data to cover a wide range of multimodal challenges and visual components. (ii) Verifiability and structured format. All questions are reformulated into a structured, open-ended QA format to facilitate objective evaluation. We exclude examples that cannot be reliably verified, such as those with incorrect answers, ambiguous phrasing, or poor readability. (iii) Appreciate difficulty. We exclude examples that the base model can easily solve and prioritize questions that remain challenging. (iv) Beneficial integration of tools. We categorize examples based on whether tool usage leads to correct answers. Cases where model can solve correctly using additional tool calls are reserved for reinforcement learning, whereas examples that remain unsolved even with tool assistance are used for cold start.
Specially, we curate data from three major categories: perception, reasoning, and search. Besides, we also include long Chain-of-Cot (CoT) reasoning data in cold start subset. Please refer to Appendix A.1 for more details on data sources. All datasets are carefully cleaned, reformatted, and divided into subsets for cold start or reinforcement learning subsets. To ensure sufficient difficulty, we employ Qwen2.5-VL-7B (bai2025qwen2) as a baseline evaluator. For each question, model is prompted to generate 8 responses, and we retain only those instances where it answers correctly at most two times, thereby filtering out trivial cases. To further assess tool-use effectiveness, we prompt model to solve each question with tool invocation, again collecting 8 responses per instance, and categorize examples according to their success rate.
Trajectories Synthesis. We construct cold start datasets by eliciting step-by-step trajectories from models (e.g., Gemini 2.5 Pro (comanici2025gemini), GPT-4o (hurst2024gpt), and Claude Sonnet 4 (claude4)). For each prompt, these models are prompted to produce detailed reasoning traces that explicitly include tool-invocation markers (e.g., code snippets). Each declared tool call is executed, and the returned outputs are fed back to the originating model, and model continues reasoning, potentially issuing further tool calls, until it produces a final answer. The entire interaction is recorded as a single trajectory. Only trajectories with correct final answers and error-free code are retained for high-quality cold-start data.
@@ 3.4 Agentic Reinforcement Learning
After cold-start training has equipped model with basic tool-use patterns, we adopt reinforcement learning to further enhance its ability to integrate tools in dynamic environment. Unlike SFT, which relies on learning from static trajectories, agentic RL places the model in an interactive environment where it must dynamically decide when and how to invoke tools in order to solve tasks. Following DeepEyes (zheng2025deepeyes), we employ a sparse and outcome-driven reward. The overall reward consists of two components: (i) accuracy reward $R_{acc}$, which evaluates whether the final answer matches the ground truth, and (ii) format reward $R_{format}$, which penalizes outputs that violate required formats. The total reward is defined as: $R=R_{acc}+R_{format}$.
@@ 4 RealX-Bench
Model Text Search Image Search Average Perception Reasoning Search Integration Proprietary & Open-source Models GPT4o hurst2024gpt 32.3 29.9 22.5 29.4 16.7 32.0 29.3 23.0 29.4 16.7 36.3 29.3 25.8 36.5 16.7 38.7 30.5 27.5 36.5 15.3 Gemini 2.5 Pro comanici2025gemini 39.3 34.8 24.2 36.0 16.7 41.7 39.0 28.7 38.4 23.6 45.0 37.8 33.2 43.1 25.0 46.0 41.5 33.7 43.6 27.8 o3 o3 35.0 31.7 23.0 30.8 11.1 41.0 34.8 28.7 37.9 19.4 41.3 37.8 28.1 40.3 22.2 39.3 38.4 25.3 37.0 20.8 Qwen2.5-VL-7B bai2025qwen2 17.0 15.9 13.5 12.3 6.9 21.7 17.7 15.7 18.5 7.6 19.7 16.6 14.4 15.9 8.3 22.3 17.1 16.3 19.9 9.7 Qwen2.5–VL-32B bai2025qwen2 25.0 21.3 19.7 19.9 12.5 25.7 25.6 20.2 19.0 16.7 30.7 27.4 23.0 26.1 15.3 32.0 27.4 29.2 31.8 23.6 Qwen2.5–VL-72B bai2025qwen2 25.3 23.1 17.4 17.5 9.7 26.3 28.7 19.7 20.4 16.7 28.0 28.7 20.2 20.4 15.3 31.0 35.4 25.8 25.6 23.6 Grounded Reasoning Models Thyme zhang2025thyme 21.0 18.3 14.6 12.8 4.2 DeepEyes zheng2025deepeyes 19.0 19.5 14.6 12.8 9.7 Agentic Multimodal Model DeepEyesV2 28.3 19.5 22.5 28.9 18.1 $\Delta$ (vs Qwen2.5-VL-7B) +6.0 +2.4 +6.2 +10.0 +8.4 Human Performance Human 70.0 69.5 63.5 62.1 51.4
Existing multimodal benchmarks such as MME-RealWorld (zhang2024mme), SEED-Bench (li2023seedbench), and MMSearch (jiang2024mmsearch) primarily evaluate isolated capabilities, for instance, perception, retrieval, or reasoning. However, real-world multimodal understanding often demands coordination across multiple abilities. Thus, we introduce RealX-Bench, a comprehensive benchmark that evaluates the coordinated interplay of perception, search, and reason in complex real-world scenarios.

### v3

# arXiv:2511.05271v3[cs.CV] 27 Feb 2026

@@ 1 Introduction
These tool-use capabilities can be categorized into two types: (i) Operation tools: Current models cannot perform complex operations on visual or numerical data, including fine-grained image manipulations (e.g., cropping, measuring) and quantitative computations. This limits their capacity to reason about detailed visual content or solve mathematical problems. (ii) Information retrieval tools: Models cannot proactively access up-to-date external knowledge, which often leads to outdated conclusions or statements without verifiable sources. Some recent works attempt to rely on a single tool. For example, as shown in Figure 1 (b), DeepEyes Zheng et al. (2025b) uses cropping to achieve fine-grained perception, but due to the lack of information retrieval capability, DeepEyes cannot correctly determine the category based solely on its internal knowledge. In contrast, although MMSearch-R1 Wu et al. (2025) can perform search, it lacks fine-grained perception, leading to retrieval failures. A substantial gap remains between existing approaches and truly agentic multimodal models. While o3 (OpenAI, 2025) has explored “thinking with image” reasoning pattern that combines operations and search, how to realize such capabilities remains unclear.
@@
We first follow the setup of DeepEyes (Zheng et al., 2025b) and apply reinforcement learning directly on Qwen2.5-VL (Bai et al., 2025), but find that limited inherent tool-use capability prevents stable tool invocation. This highlights the need for a cold-start stage to establish reliable tool-use patterns. Thus, we curate a high-quality dataset that spans diverse scenarios, including perception, reasoning, and search tasks. After cleaning, we apply two filters: (i) difficulty filtering, retaining only questions unsolvable by the base model, and (ii) tool-benefit classification, keeping cases where tool use improves accuracy. Data are split into two subsets: tool-solvable examples for RL and harder unsolved cases for cold start, further augmented with long chain-of-thought trajectories. Supervised fine-tuning on this cold-start dataset enables the model to acquire basic tool-use patterns and deeper reasoning, after which RL further strengthens tool invocation. Notably, we rely only on two simple rewards, accuracy and format, without complex reward engineering (Su et al., 2025).
@@
Besides, we evaluate DeepEyesV2 on benchmarks covering real-world understanding, mathematical reasoning, and search-intensive tasks. DeepEyesV2 outperforms both general-purpose MLLMs and prior specific reasoning approaches. Specifically, on real-world understanding benchmarks, DeepEyesV2 surpasses even Qwen2.5-VL-32B in some benchmarks through effective tool use. On reasoning tasks, DeepEyesV2 shows preformance gains across multiple benchmarks, including +7.1+7.1 on MathVerse ($52.7\%$ accuracy). On search benchmarks, DeepEyesV2 delivers strong advantages, reaching $63.7\%$ on MMSearch (Jiang et al., 2024), far beyond the MMSearch-R1 (Wu et al., 2025) ($53.8\%$). These results demonstrate that by reliably invoking tools, DeepEyesV2 extends its comprehensive capabilities, achieving accurate and advanced reasoning.
@@ 2 Related Works
Multimodal Large Language Models. The field of multimodal large language models (MLLMs) has witnessed rapid progress in recent years. Early efforts mainly focus on combining pretrained visual encoders with large language models through lightweight adapters or projection layers, enabling basic vision–language alignment and simple multimodal understanding (Li et al., 2023a; Liu et al., 2023, 2024a; Bai et al., 2023; Chen et al., 2024). Subsequently, more powerful architectures such as Qwen2.5-VL (Bai et al., 2025), LLaVA-OneVision (Li et al., 2024a), and InternVL3 (Zhu et al., 2025), expand the training scale and integrated more diverse visual data, significantly improving performance on benchmarks of visual question answering, captioning, and general perception tasks. Recently, some OmniMLLMs (Li et al., 2025b; Zhao et al., 2025a; Fu et al., 2024; Jain et al., 2024; Hong et al., 2025) are capable of processing a mix of modalities like speech, video, and images simultaneously. However, existing MLLMs remain largely passive: they can interpret multimodal inputs and generate answers, but lack the ability to actively invoke external tools for computation or knowledge retrieval, which limits their reliability in complex reasoning tasks.
Thinking with Images. The paradigm of “Think with Image” is first introduced by o3 (OpenAI, 2025), which demonstrated that multimodal models can interleave reasoning with iterative visual analysis, actively manipulating images to support step-by-step problem solving. Many works attempt to reproduce such capabilities. Most approaches (Su et al., 2025; Lai et al., 2025; Liu et al., 2025; Fan et al., 2025; Jiang et al., 2025; Zhang et al., 2025a) adopt a two-stage training pipeline, where a cold-start phase is followed by reinforcement learning. In contrast, DeepEyes (Zheng et al., 2025b) only adopts reinforcement learning alone and incentivizes the “Think with Image” behaviors, leading to strong reasoning performance. However, the majority of these efforts employ a rather limited tool set, typically restricted to region cropping for fine-grained perception. To improve generality, PyVision (Zhao et al., 2025b) and Thyme (Zhang et al., 2025b) utilize code execution to enable more flexible visual operations. Despite this progress, these models remain constrained to image manipulation only, and are unable to handle knowledge-intensive questions where access to up-to-date external information is essential.
@@
Search-oriented Reasoning. To mitigate the inherent knowledge limitations of large multimodal language models, a growing line of work explores augmenting them with external knowledge acquisition. Early approaches commonly adopt the retrieval-augmented generation (RAG) paradigm (Song et al., 2025; Jin et al., 2025), where relevant information is retrieved from a pre-constructed knowledge base and fed into the model. While effective, this paradigm remains constrained by the static and finite nature of the underlying corpus. To overcome these limitations, more recent studies attempt to leverage online search to dynamically access broader and up-to-date information (Zheng et al., 2025a). Beyond purely textual queries, some efforts extend search into the multimodal domain, enabling retrieval of not only documents but also images, charts, or other media forms relevant to the task (Wu et al., 2025; Wang et al., 2025d). These advances highlight the potential of search-augmented reasoning to complement perception and tool-use capabilities, ultimately broadening the scope of problems that multimodal models can effectively address.
@@ 3.1 Overall Pipeline
Similar to DeepEyes (Zheng et al., 2025b), DeepEyesV2 is an agentic multimodal model, but with extended tool-use capabilities beyond simple cropping. In DeepEyesV2, programmatic code execution and web retrieval are treated as complementary and interleavable tools inside a single reasoning trajectory (see Figure 3). Given an image input and the corresponding user query, DeepEyesV2 first generates an initial reasoning plan, and explicitly determines whether this question can be solved directly through internal reasoning or requires tool invocation. If tool use is necessary, DeepEyesV2 emits executable Python code or issues web search queries. Code execution is carried out in a sandboxed environment and can produce structured outputs such as transformed images, numerical measurements, computed arrays, plots, or execution logs. Image queries using the original whole image are submitted via SerpAPI and return the top five visually matched webpages (each with a thumbnail and title). Text queries return the five most relevant webpages, along with titles and snippets. All tool outputs are converted into observations and appended to model’s context. DeepEyesV2 then thinks further in light of these observations and may plan further tool invocations (either additional code, further searches, or both), iterating this reasoning–tool–integration loop until a conclusive answer is produced.
@@ 3.2 Pioneer Experiments
To investigate whether MLLMs can directly acquire tool-use ability through reinforcement learning, we first conduct a pioneer experiment on Qwen2.5-VL (Bai et al., 2025) following DeepEyes (Zheng et al., 2025b). As shown in Figure 4, during training, we observe that in the early stages model occasionally attempts to produce Python code, but these outputs are often buggy or fail to execute, indicating that existing MLLMs struggle to generate stable and reliable code. As training continues, model gradually abandons code generation and converges to producing only short reasoning chains followed by direct answers, thereby bypassing tool use. Then, to encourage tool invocation, we incorporate the tool usage bonus mechanism from DeepEyes, which explicitly rewards the generation of code. With this additional signal, model is indeed able to produce correct and runnable code in the early stages, suggesting that the mechanism can enforce coding ability. However, with continued training a new degeneration emerges: model’s behavior converged to emitting exactly one code block per query, and this single block typically consists of non-executable, placeholder comments rather than meaningful code, revealing the phenomenon of reward hacking. This pioneer experiment highlights that existing MLLMs cannot reliably learn complex tool use through direct RL alone, motivating the need for a cold start to bootstrap model’s tool invocation ability.
@@ 3.3 Training Data Curation
Data Collection. Pioneer experiments have highlighted the necessity of constructing a high-quality dataset for supervised fine-tuning to explicitly guide model to learn how to generate executable code and perform tool invocations. Following DeepEyes (Zheng et al., 2025b), we collect data in accordance with the following principles: (i) Diverse tasks and image distribution. We incorporate varied data to cover a wide range of multimodal challenges and visual components. (ii) Verifiability and structured format. All questions are reformulated into a structured, open-ended QA format to facilitate objective evaluation. We exclude examples that cannot be reliably verified, such as those with incorrect answers, ambiguous phrasing, or poor readability. (iii) Appreciate difficulty. We exclude examples that the base model can easily solve and prioritize questions that remain challenging. (iv) Beneficial integration of tools. We categorize examples based on whether tool usage leads to correct answers. Cases where model can solve correctly using additional tool calls are reserved for reinforcement learning, whereas examples that remain unsolved even with tool assistance are used for cold start.
Specially, we curate data from three major categories: perception, reasoning, and search. Besides, we also include long Chain-of-Cot (CoT) reasoning data in cold start subset. Please refer to Appendix A.1 for more details on data sources. All datasets are carefully cleaned, reformatted, and divided into subsets for cold start or reinforcement learning subsets. To ensure sufficient difficulty, we employ Qwen2.5-VL-7B (Bai et al., 2025) as a baseline evaluator. For each question, model is prompted to generate 8 responses, and we retain only those instances where it answers correctly at most two times, thereby filtering out trivial cases. To further assess tool-use effectiveness, we prompt model to solve each question with tool invocation, again collecting 8 responses per instance, and categorize examples according to their success rate.
Trajectories Synthesis. We construct cold start datasets by eliciting step-by-step trajectories from models (e.g., Gemini 2.5 Pro (Comanici et al., 2025), GPT-4o (Hurst et al., 2024), and Claude Sonnet 4 (Anthropic, 2025)). For each prompt, these models are prompted to produce detailed reasoning traces that explicitly include tool-invocation markers (e.g., code snippets). Each declared tool call is executed, and the returned outputs are fed back to the originating model, and model continues reasoning, potentially issuing further tool calls, until it produces a final answer. The entire interaction is recorded as a single trajectory. Only trajectories with correct final answers and error-free code are retained for high-quality cold-start data.
@@ 3.4 Agentic Reinforcement Learning
After cold-start training has equipped model with basic tool-use patterns, we adopt reinforcement learning to further enhance its ability to integrate tools in dynamic environment. Unlike SFT, which relies on learning from static trajectories, agentic RL places the model in an interactive environment where it must dynamically decide when and how to invoke tools in order to solve tasks. Following DeepEyes (Zheng et al., 2025b), we employ a sparse and outcome-driven reward. The overall reward consists of two components: (i) accuracy reward $R_{acc}$, which evaluates whether the final answer matches the ground truth, and (ii) format reward $R_{format}$, which penalizes outputs that violate required formats. The total reward is defined as: $R=R_{acc}+R_{format}$.
@@ 4 RealX-Bench
Model Text Search Image Search Average Perception Reasoning Search Integration Proprietary & Open-source Models GPT4o Hurst et al. (2024) 32.3 29.9 22.5 29.4 16.7 32.0 29.3 23.0 29.4 16.7 36.3 29.3 25.8 36.5 16.7 38.7 30.5 27.5 36.5 15.3 Gemini 2.5 Pro Comanici et al. (2025) 39.3 34.8 24.2 36.0 16.7 41.7 39.0 28.7 38.4 23.6 45.0 37.8 33.2 43.1 25.0 46.0 41.5 33.7 43.6 27.8 o3 OpenAI (2025) 35.0 31.7 23.0 30.8 11.1 41.0 34.8 28.7 37.9 19.4 41.3 37.8 28.1 40.3 22.2 39.3 38.4 25.3 37.0 20.8 Qwen2.5-VL-7B Bai et al. (2025) 17.0 15.9 13.5 12.3 6.9 21.7 17.7 15.7 18.5 7.6 19.7 16.6 14.4 15.9 8.3 22.3 17.1 16.3 19.9 9.7 Qwen2.5–VL-32B Bai et al. (2025) 25.0 21.3 19.7 19.9 12.5 25.7 25.6 20.2 19.0 16.7 30.7 27.4 23.0 26.1 15.3 32.0 27.4 29.2 31.8 23.6 Qwen2.5–VL-72B Bai et al. (2025) 25.3 23.1 17.4 17.5 9.7 26.3 28.7 19.7 20.4 16.7 28.0 28.7 20.2 20.4 15.3 31.0 35.4 25.8 25.6 23.6 Grounded Reasoning Models Thyme Zhang et al. (2025b) 21.0 18.3 14.6 12.8 4.2 DeepEyes Zheng et al. (2025b) 19.0 19.5 14.6 12.8 9.7 Agentic Multimodal Model DeepEyesV2 28.3 19.5 22.5 28.9 18.1 $\Delta$ (vs Qwen2.5-VL-7B) +6.0 +2.4 +6.2 +10.0 +8.4 Human Performance Human 70.0 69.5 63.5 62.1 51.4
Existing multimodal benchmarks such as MME-RealWorld (Zhang et al., 2024b), SEED-Bench (Li et al., 2023b), and MMSearch (Jiang et al., 2024) primarily evaluate isolated capabilities, for instance, perception, retrieval, or reasoning. However, real-world multimodal understanding often demands coordination across multiple abilities. Thus, we introduce RealX-Bench, a comprehensive benchmark that evaluates the coordinated interplay of perception, search, and reason in complex real-world scenarios.